OpenStudy (anonymous):

Solve by factoring x^2+3x=8x-6

OpenStudy (anonymous):

x^2 + 3x = 8x - 6 x^2 = 8x-3x - 6 x^2= 5x - 6 x^2-5x + 6 = 0 x^2 - 3x- 2x + 6= 0 x(x-3) -2(x-3) (x-2)(x-3) = 0 x=2 or x= 3